OpenStudy (commonlizer):

How many solutions does the equation 5m − 5m − 12 = 14 − 2 have? Zero One Two Many

OpenStudy (commonlizer):

these Questions are confusing to me plz hlp

OpenStudy (anonymous):

5m – 5m – 12 = 14 – 2 (5 - 5)m - 12 = -12 0m = -12 + 12 0m = 0 Zero times any number is zero. Simplifying 5m + -5m + -12 = 14 + -2 Reorder the terms: -12 + 5m + -5m = 14 + -2 Combine like terms: 5m + -5m = 0 -12 + 0 = 14 + -2 -12 = 14 + -2 Combine like terms: 14 + -2 = 12 -12 = 12 Solving -12 = 12 Couldn't find a variable to solve for. This equation is invalid, the left and right sides are not equal, therefore there is no solution.

OpenStudy (anonymous):

Either those two

OpenStudy (anonymous):

5m − 5m − 12 = 14 − 2 0 - 12 = 14 - 2 -12 = 12 Or This one hope thes comment I made helps u